### v1

# SpecEval: Evaluating Code Comprehension in Large Language Models via Program Specifications

Abstract.
Large Language models (i.e. LLMs) have achieved impressive performance in automated software engineering. Extensive efforts have been made to evaluate the abilities of code LLMs in various aspects, with an increasing number of benchmarks and evaluation frameworks proposed. Apart from the most sought-after capability of code generation, the capability of code comprehension is being granted growing attention. Nevertheless, existing works assessing the code comprehension capability of LLMs exhibit varied limitations. For example, evaluation frameworks like CRUXEval and REval usually focus on code reasoning tasks over a certain input case. It leads to a limited range of execution traces covered, resulting in a loss in code semantics examined and the inability to assess the comprehensive understanding of LLMs concerning the target program.
To tackle the aforementioned challenges, we propose SpecEval, a novel black-box evaluation framework to evaluate code comprehension in LLMs via program specifications. Inspired by the idea that specifications can act as a comprehensive articulation of program behaviors concerning all possible execution traces, we employ formalized program specifications to represent program semantics and perform comprehensive evaluations. In particular, four specification-related tasks are designed meticulously to assess the capability of LLMs from basic to advanced levels. Moreover, counterfactual analysis is conducted to study the performance variance of LLMs under semantics-preserving perturbations, and progressive consistency analysis is performed to study the performance consistency of LLMs over a series of tasks with sequential dependence. Systematic experiments are conducted on six state-of-the-art LLMs. Extensive experimental results present a below-satisfactory performance of LLMs on specification-related tasks, revealing the limitations of existing LLMs in terms of articulating program semantics. The counterfactual analysis and progress consistency analysis also reveal the sensitivity of LLMs towards semantic-preserving perturbations and limited progressive consistency across the progressive tasks. Our experimental results and analysis underscore future directions for enhancement.
Large language models, Program Semantics, Program Specifications
ccs: General and reference Empirical studiesccs: Theory of computation Program specifications
1. Introduction
Automated software engineering, encompassing areas such as code generation (Liu et al., 2024b; Chen et al., 2021), program repair (Xia et al., 2023; Nilizadeh et al., 2021), and code summarization (Zhang et al., 2020; McBurney and McMillan, 2015), has persistently remained a prominent research focus in both academic and industrial practice. Early research in this field adopted traditional techniques, such as template-based (Flanagan and Leino, 2001; Ernst et al., 2007) and retrieval-based (Hayati et al., 2018; Zhang et al., 2020) methods, to address these issues. However, due to technology and data availability constraints, these methods often struggled to deliver satisfactory performance. With the rise and advancement of deep learning technologies, an increasing number (Yu et al., 2019), Transformer (Vaswani, 2017), and GNNs (Wu et al., 2020), have been applied to this field. Extensive research on program semantics learning with advanced neural network models drives significant progress and advancements in various SE applications. Program semantics are the essence of a program, programs with identical semantics may have different expressions. Consequently, numerous studies (Susanto and Lu, 2017; Zhou et al., 2019; Peng et al., 2021) have explored different techniques, including tree-based and graph-based approaches, to improve the neural networks in learning program semantics.
Recently, large language models (LLMs) (OpenAI, 2023; Achiam et al., 2023; Wei et al., 2023; Nijkamp et al., 2023; Lozhkov et al., 2024; Guo et al., 2024; Touvron et al., 2023; Roziere et al., 2023) have impressed the community with their outstanding performance, vastly surpassing previous techniques in various software development applications. In the meantime, a wide range of software development tools (Github, 2024; JeyBrains, 2024) powered by these models have significantly increased the productivity of software developers. Behind the rapid advancement of these LLMs, compared to earlier techniques such as graph-based approaches that explicitly incorporated code semantics into model training to help the model comprehend program semantics, LLMs usually treat the code snippet as pure text and rely on the straightforward pre-training techniques i.e., next-token prediction to train on a massive amount of code corpus. As a result, it raises a question i.e., does the simple next-token pertaining technique truly enable models to comprehend the complex semantics of code?
Numerous benchmarks, such as HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and SWE-bench (Jimenez et al., 2023), are proposed to evaluate the coding capacity of LLMs. Although these benchmarks can, to some extent, reflect the code generation capabilities of different models, using the criterion of whether the generated code passes test cases is still insufficient to determine whether the models genuinely understand the code (Yang et al., 2024). To conduct a more in-depth analysis, CRUXEval (Gu et al., 2024) and REval (Chen et al., 2024) define certain program execution tasks, such as code coverage prediction and output prediction, to assess the code reasoning capabilities of LLMs by comparing the model’s generated results with actual program runtime outcomes. Nevertheless, the dynamic execution results of a program depend on variations in its input, making it challenging to ensure that a comprehensive set of inputs can cover all possible execution traces  [^1] The execution trace refers to the transitions of program states (Plotkin, 1981) where the states are defined as the set of all program variables including their types and values.. It leads to limited program inputs invoking limited execution traces for a program and cannot comprehensively reflect the full program semantics. Some probing techniques (Hernández López et al., 2022; Wan et al., 2022; Ma et al., ) are adopted to analyze the code understanding, such as code syntax and semantic knowledge. They typically require access to a model’s intermediate vector representations for corresponding inputs to investigate whether the model has learned specific attributes. As these techniques usually require direct access to the model, they are not feasible for black-box models such as GPT-3.5 and GPT-4.
These challenges inspire us to explore other forms of representation for program semantics, which are required to articulate program behaviors in multiple granularities. Formalized program specifications encompass precise claims that describe the intended or actual program behaviors in formal languages. By specifying the constraints on program variables or the transition of program states, the strong program specifications can effectively articulate program semantics, accurately capturing program behaviors either generally or in detail. Moreover, the correctness of formal specifications can be effectively validated by corresponding specification verifiers (Cok, 2011; Ahrendt et al., 2014; Kirchner et al., 2015). It also overcomes the limitations of using natural languages, such as code summaries, to describe the functionality of code, which cannot be automatically verified.
Inspired by these ideas, we propose SpecEval, a novel black-box evaluation framework featuring the representation of program semantics in the form of formal program specifications. We define multiple specification-related tasks for subject LLMs to complete. These tasks require LLMs to demonstrate an understanding of code by summarizing the code behavior and generating formalized statements. By leveraging automated procedures for determining the correctness of specifications, the assessment can be conducted rigorously and efficiently. In particular, four tasks with sequential dependencies are designed, including Judgement, Selection, Infilling, and Generation. In the Judgement task, LLMs are required to evaluate the correctness of the given specifications for the code. The Selection task involves selecting the most appropriate candidate specification for the target program. The Infilling task requires LLMs to complete partially provided specifications by filling in the correct sub-expressions. Finally, the Generation task challenges LLMs to generate correct specifications entirely from scratch. For each task, we further introduce counterfactual analysis to cast code comprehension as the problem of determining how controlled input code changes result in model output changes by the semantic-preserving perturbations, with mutation operators including Def-use Break, If-else Flip, Independent Swap, Name Random, and Name Shuffle. Moreover, for each kind of perturbation, Progressive Consistency Analysis is conducted with the well-designed metric Progressive Consistency Score to examine the consistency performance of LLMs over these tasks with progressive dependencies, enabling the framework to assess the code comprehension capacities of LLMs from the shallow to the deep.
To construct the framework, we collect 204 Java programs with verifiable ground-truth specifications in JML (Leavens et al., 1998) style from existing datasets, including Frama-C-Problems (github, 2024), the Java category of SV-COMP Benchmark (sosy lab, 2024), and a dataset collected by Ma et al.(Ma et al., 2024). Test cases for these programs are also generated to assist the specification validation process. We conduct experiments on six large language models, including state-of-the-art code LLMs and general LLMs, for a systematic comparison between the popular models. Experiment results show that LLMs still have a long way to go before they can fully articulate program semantics with formal specifications, with less than satisfactory performance in most specification-related tasks. Also, Counterfactual Analysis discloses the sensitivity of LLMs to different extents towards semantic-preserving perturbations. Additionally, Progressive Consistency Analysis suggests that the reasoning ability of LLMs does not work in a progressive manner as humans do.
In summary, the main contributions of our work include:
  • A novel black-box evaluation framework assessing the program semantics learning abilities of LLMs, featuring the representation of program semantics in the form of formalized specifications. Four tasks with progressive difficulties and Counterfactual analysis are designed to study the performance of LLMs from multiple aspects.
  • An adapted benchmark (Author, s) based on existing datasets (sosy lab, 2024; github, 2024; Nilizadeh et al., 2021; Ma et al., 2024), consisting of 204 Java programs, along with ground-truth specifications, test cases for dynamic validation, and five semantics-equivalent variants, is released to facilitate follow-up research.
  • A thorough evaluation of six state-of-the-art code LLMs and general LLMs, revealing the limitations of LLMs concerning the expression of program semantics with formal specifications, underscoring the directions for future enhancement of LLMs.
2. Motivation and Related work
2.1. Program Specification
Program specification refers to precise statements that define a program’s intended or actual behaviors, either in its entirety or in specific components. According to the language adopted, program specifications can be categorized into natural language specifications and formalized specifications. Natural language specifications primarily encompass software documentation and code comments. Another significant portion of program specifications is articulated using formal languages, such as mathematical expressions (Leavens et al., 1998; Kirchner et al., 2015) to specify the constraints governing a program’s behavior, and automata (Lo and Khoo, 2006; Shoham et al., 2007) to describe the transitions of program states. Compared with natural language, formalized specifications are rigorous and unambiguous. Program specifications have been widely adopted in a variety of software engineering tasks, such as requirement engineering (Aman and Ibrahim, 2013; Sharifi et al., 2020), software testing (Mesbah et al., 2012; Nguyen and Dang, 2023), and model checking (Cabodi et al., 2008; Beyer et al., 2015).
In SpecEval, we mainly focus on formal specifications written in Java Modeling Language (JML) (Leavens et al., 1998) describing the actual behaviors of Java programs. One example program isPalindrome and corresponding ground-truth specifications are illustrated in Fig. 1. The program aims to check whether a given input string s is a palindrome string, i.e., a string that reads the same forward and backward. The specifications are instrumented within the program in the form of comments starting with //@. Generally, the specifications involved can be divided into three categories:
  • Preconditions (line 1), described in requires statements, specifying the constraints on input variables that must be met before the corresponding code (i.e. isPalindrome here) is executed.
  • Postconditions (line 2), described in ensures statements, specifying the properties of the returned value that always hold after the corresponding code is executed.
  • Loop Invariants (lines 5, 6, and 7), described in maintaining statements, specifying the constraints on the local variables that always hold each time before the loop body is executed.
Together, the specifications mentioned above constitute a detailed articulation of the semantics of the target program. For the method isPalindrome as a whole, preconditions and postconditions specify its overall requirements and functionalities. For the internal implementation of isPalindrome, loop invariants reveal the pattern followed by local variables when their values are updated.
Figure 1. An example program and corresponding ground-truth JML specification (highlighted in blue).
It is worth noting that formalized program specifications are more rigorous, eliminating ambiguity compared with specifications in natural language. Moreover, formalized program specifications can thoroughly articulate program semantics. Program semantics essentially lies in the evolution of program states, which are usually defined by the set of variable values. By specifying the constraints on program variables or the transition of program states, the strong program specifications are rich in semantic information and can accurately capture program behaviors. In addition, the correctness of formal specifications can be effectively validated. A good number of specification verifiers (Cok, 2011; Ahrendt et al., 2014; Kirchner et al., 2015) is proposed to automatically check the consistency between the given specifications and the target program. Consequently, compared with natural language specifications, the correctness of program specifications can be automatically verified.
2.2. Program Semantics Comprehension of LLMs
Program semantics are the core of a program, and only when code models can effectively learn these semantics can they provide reliable support and assurance in software development. Recently, there has been a surge in the impact of LLMs, playing a significant role in various software development scenarios. In light of the critical role of code semantics, a substantial body of research has attempted to evaluate whether LLMs genuinely comprehend code through various analytical techniques.
A typical technique involves using probing analysis (Hernández López et al., 2022; Wan et al., 2022; Ma et al., ) to explore whether code models have learned specific code attributes from the trained code datasets. Specifically, it leverages the transformation of code semantic information, such as data flow, into high-dimensional vector representations in the latent space. This enables an examination of whether the model truly understands the semantics of the code. Yet, they are usually required to access the model’s intermediate vector representations for analysis, which is not feasible for black-box commercial models such as GPT-3.5 and GPT-4. Moreover, some popular code generation benchmarks such as HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and SWE-bench (Jimenez et al., 2023) are proposed to evaluate the coding capacity of LLMs. Yet, the criterion to judge whether the generated code can pass test cases is insufficient to articulate the model’s understanding of the code.
The latest research REval (Chen et al., 2024) proposed a framework to evaluate the code reasoning ability of LLMs by inferring the program runtime behaviors. Inspired by CRUXEval (Gu et al., 2024), REval proposed four evaluation tasks, including Code Coverage Prediction, Program State Prediction, Execution Path Prediction, and Output Prediction for evaluation. In particular, Code Coverage Prediction requires LLMs to predict whether a statement will be executed given a certain input. Program State Prediction requires LLMs to reason about the type and value of a specific variable. Execution Path Prediction requires LLMs to predict the next statement executed under certain circumstances, and Output Prediction requires LLMs to generate the program’s output directly from the given input.
It should be emphasized that in each of the four tasks mentioned, LLMs are provided with one specific input respectively. For a deterministic program, one specific input can only result in one particular trace of program state transition, i.e. the execution trace. This observation leads to the fact that for each independent task assigned, LLMs are only reasoning about program behaviors within one specific execution trace, whereas there exists an infinite number of execution traces (corresponding to the infinite number of inputs) within the program state space other than the one being examined. Nevertheless, program semantics is not only about the program behaviors during one specific execution but also about the general pattern of program state transition adaptable to all potential program states. The set of program behaviors for all potential program states and execution traces constitutes full program semantics. This indicates that the practice employed by REval could cause a loss in the program semantics examined by its tasks, since REval can only supply a limited number of evaluation tasks and inputs, leaving the other program behaviors in an infinite number of execution traces uncovered.
[…]
3. Methodology
In this section, we aim to present a detailed introduction to our framework, SpecEval. We first provide an overview of SpecEval, then describe the evaluation tasks in our framework. For each task, we further introduce the counterfactual analysis, framing code comprehension as the challenge of assessing how controlled input code modifications, through semantic-preserving perturbations, lead to changes in the model’s output. Moreover, for each type of perturbation, we perform Progressive Consistency Analysis to evaluate the consistency of LLM performance across tasks of increasing difficulty. This framework enables a comprehensive assessment of the code comprehension capabilities of LLMs, ranging from superficial to deeper levels of understanding.
3.1. Overview
Fig. 2 demonstrates the overview of our framework, which is characterized by employing formalized program specifications to represent program semantics. We generally adopt the typical architecture of existing black-box evaluations (Chen et al., 2024; Gu et al., 2024), where several evaluation tasks are assigned to the LLMs. The ability of LLMs is evaluated by assessing their performance in these tasks. In our frameworks, four tasks related to program specifications are designed, including Specification Correctness Judgement, Specification Candidates Selection, Specification Infilling, and Specification Generation. For each task, we design five semantic-preserving perturbations performed on the programs and the corresponding ground-truth specifications to create variants with equivalent semantics. Counterfactual analysis is conducted based on the original and mutants, allowing us to study the performance variance of LLMs against semantic-preserving perturbations. Progressive consistency analysis is performed for each type of perturbation to evaluate the consistency of LLM performance on a series of tasks with sequential dependencies in between.
Figure 2. Overview of the evaluation framework SpecEval.
3.2. Evaluation Tasks and Metrics
We designed four tasks related to formal program specifications. These tasks not only request LLMs to summarize abstract semantic information into formalized statements but also require them to comprehend the given specifications and take appropriate actions based on their understanding.
3.2.1. Specification Correctness Judgement
Specifications possess correctness, defined by their consistency with program behaviors. A correct specification is about constraints that always hold for all program states in its scope. We denote the actual correctness of the candidate specification $s$ as a function $\mathbf{Cr}:S\rightarrow\{\mathbf{true},\mathbf{false}\}$, which takes a value of $\mathbf{Cr}(s)=\mathbf{true}$ if $s$ is correct and $\mathbf{Cr}(s)=\mathbf{false}$ otherwise. In this task, subject LLM is exploited to decide the correctness of a given candidate specification on a target program. The correct candidates are extracted from the ground-truth specifications of the target program, and the incorrect candidates are obtained by modifying essential components within the ground-truth specification, including replacing variable names with other variables in the same scope, altering variable operators, and swapping statement predicates (i.e., between $\forall$ and $\exists$). Each component of interest is decided randomly whether it will be modified, with equal probabilities (at 50%) of being modified or staying unchanged. At least one modification is performed to ensure the specification is not exactly the same.
  • Task Formalism. Given a target program $P$, and a candidate specification $s$, a judgment task $t\in T$ is a 2-tuple $(P,s)$, where the subject model $\mathcal{M}:T\rightarrow\{\mathbf{true},\mathbf{false}\}$ is required to decide whether the given specification is correct for the target program.
  • Evaluation Metrics. In this task, Accuracy measures the percentage of correctly judged specifications. The Accuracy for this task can be calculated by the following formula:
    $Accuracy=\frac{1}{|T|}\sum\limits_{(P,\ s)\in T}\mathbb{I}(\mathcal{M}(P,s)=\mathbf{Cr}(s))$
    where $\mathbb{I}:\{\mathbf{true},\mathbf{false}\}\rightarrow\{0,1\}$ is the indicator function that returns 1 if and only if the given boolean expression takes a value of $true$.
3.2.2. Specification Candidates Selection
Specifications possess varying degrees of strength. While weaker specifications such as ensures true can still pass the verification and should be considered correct, they often describe only trivial information about the program. In contrast, stronger specifications provide a more comprehensive expression of the program’s semantics. Therefore, in this task, we consider both the correctness and strength when constructing candidates for the given target program. In particular, four candidate specifications are provided to subject LLM, which is required to choose the most appropriate (i.e., correct and strongest) one for the given target program. The most appropriate candidate (denoted as $\hat{s}$) in four candidates is extracted from the ground-truth specifications for the target program. Other candidates have an equal chance of being modified specifications or trivial specifications. The former is obtained by modifying the ground-truth specification with the same logic mentioned in Section 3.2.1. The latter is randomly chosen from a set of pre-defined simple specifications that are too weak to be incorrect, such as ensures true.
  • Task Formalism. Given a target program $P$, and a set of four candidate specifications $S=\{s_{1},s_{2},s_{3},s_{4}\}$, a selection task $t\in T$ is a 2-tuple $(P,S)$, where the subject model $\mathcal{M}:T\rightarrow S$ is required to return the most appropriate candidate within $S$ for the target program.
  • Evaluation Metrics. In this task, Accuracy measures the percentage of correctly selected specifications. The Accuracy for this task can be calculated by the following formula.
    $Accuracy=\frac{1}{|T|}\sum\limits_{(P,\ S)\in T}\mathbb{I}(\mathcal{M}(P,S)=\hat{s})$
3.2.3. Specification Infilling
In this task, an uncompleted specification (partially masked with a placeholder) is provided to the subject LLM, which is required to fill in the placeholder with appropriate sub-expressions according to the given target program so that the final specification can pass the verifier. The uncompleted specification is acquired by randomly masking specific components of the ground-truth specifications with placeholders. At most, one placeholder is inserted once. The components of interest include array index, variable and method names, and boundary constraints of $\forall$ and $\exists$ statements, which can be identified by parsing specifications into Abstract Syntax Trees.
  • Task Formalism. Given a target program $P$, and an uncompleted specification $s$, an infilling task $t\in T$ is a 2-tuple $(P,s)$, where the subject model $\mathcal{M}:T\rightarrow S$ is required to fill in the placeholder of $s$ according to the target program and return the filled-in version of $s$.
  • Evaluation Metrics. In this task, we use #Pass and Accuracy for measurement. #Pass denotes the number of programs $P$ for which the specification infilled by model $\mathcal{M}$ are correct. Accuracy measures the percentage of correctly infilled specifications. Denoting the correctness of specification $s$ as $\mathbf{Cr}(s)$, the Accuracy for this task can be calculated as:
    $Accuracy=\frac{1}{|T|}\sum\limits_{(P,\ s)\in T}\mathbb{I}(\mathbf{Cr}(\mathcal{M}(P,s)))$
3.2.4. Specification Generation
In this task, the subject LLM is provided with a program $P$ with no instrumented specifications. The LLM is exploited to generate a set of specifications $S$ from scratch, describing the general behaviors and functionalities of the target program as closely as possible. The generated specifications are required to include pre/post-conditions for all methods and loop invariants for all loops involved in the target program.
  • Task Formalism. Given a target program $P\in\mathbb{P}$, the subject model $\mathcal{M}:\mathbb{P}\rightarrow 2^{S}$ is required to generate an appropriate set of specifications $S$ for $P$.
  • Evaluation Metrics. In this task, Precision and Recall are adopted to measure the performance of LLMs. Precision measures the percentage of correct specifications that can pass the verifier over all the generated specifications for the target program $P$. Recall measures the percentage of correctly generated ground-truth specifications (i.e., correct and strongest) over all the ground-truth specifications for the target program $P$.
    $Precision(P)=\frac{\sum_{s\in\mathcal{M}(P)}\mathbb{I}(\mathbf{Cr}(s))}{|\mathcal{M}(P)|}\quad Recall(P)=\frac{\sum_{s\in\mathcal{M}(P)}\mathbb{I}(s\in\hat{S}_{P})}{|\hat{S}_{P}|}$
    where $\hat{S}_{P}$ denotes the set of all ground-truth specifications for program $P$. Correspondingly, the average precision and recall on the set $\mathbb{P}$ of all programs in the benchmark can be calculated as:
    $AvgPrecision=\frac{1}{|\mathbb{P}|}\sum\limits_{P\in\mathbb{P}}Precision(P)\quad AvgRecall=\frac{1}{|\mathbb{P}|}\sum\limits_{P\in\mathbb{P}}Recall(P)$
    Apart from precision and recall, we also use #Pass metric to denote the number of programs $P$ for which all specifications generated by model $\mathcal{M}$ can pass the verifier.
3.3. Counterfactual Analysis
Counterfactual Analysis is a specific technique conducted by observing the variance in the model performance after changing the model inputs in a particular way (Verma et al., 2020). In SpecEval, for each kind of the designed evaluation task in Section 3.2, we respectively adopt some typical semantic-preserving perturbations to analyze the performance of LLMs on the task.
3.3.1. Semantic-preserving Perturbations
To eliminate potential bias, the perturbations should be performed while adhering to the following criteria.
  • Consistency with the original program. The perturbations should not alter the overall behaviors and functionalities of the program. The perturbed program and the original program should perform exactly the same behaviors if treated as black boxes.
  • Consistency with the original specifications. All specifications for the original program should also be adaptable to the perturbed program without significant structural or semantic modifications. Trivial modifications that have no semantic effects are allowed, such as altering the variable names involved according to specific patterns. The perturbation should not reduce the program points of interest (e.g. methods and loops) where specifications should be generated.
  • Consistency with the evaluation tasks. The perturbed programs should still be adaptable to all the tasks designed for the original program. The perturbations should not alter the problems or the corresponding answers (i.e., ground truth) involved in the tasks. Similar to the previous rule, trivial modifications such as the variable name modification can be performed on the prompts engaged in the tasks if necessary.
Figure 3. Illustrations for the semantic-preserving perturbations, including the original program, corresponding specifications, and all perturbed programs. The code and specifications modified by the perturbations are highlighted in blue.
Following prior research (Hooda et al., 2024), we adopted five different types of semantic-preserving perturbations that meet the criteria above, including Def-use Break, If-else Flip, Independent Swap, Name random, and Name shuffle.
  • Def-use Break. Def-use chain refers to the relationship between the definitions of a certain variable (where it is assigned with a value, e.g., x = 0) and its subsequent uses (where its value is accessed, e.g., func(x)). The Def-use Break perturbation aims to break the def-use chains within target programs. This is done by assigning the value of a variable var1 to a newly introduced variable var2 and altering all subsequent uses of var1 to var2. In our implementations, the variable name for var2 consists of five random letters or digits. For instance, in the program shown in Fig. 3, a new variable PZYcz inherits the value of variable num1, and replaces the latter in the remaining part of the program. Similar modifications are performed on variable num2 and local variable temp. Since the newly introduced variable completely supersedes the role of the original variable, the semantic-preserving feature of this perturbation is guaranteed.
  • If-else Flip. This perturbation swaps the branches of an if-else statement. This is done by replacing the content of the two branches with each other and negating the branch condition. Obviously, such perturbation does not affect the semantics of the original program.
  • Independent Swap. For ease of handling, we adopt a safe definition for independent statements. For two adjacent statements $S_{1}$ and $S_{2}$ in the same basic block, denote the sets of variables defined and used by statement $S_{i}$ as $V^{def}_{i}$ and $V^{use}_{i}$ respectively. If the following three conditions are satisfied simultaneously: (1) $V^{def}_{1}\cap V^{def}_{2}=\emptyset$ (2) $V^{use}_{1}\cap V^{def}_{2}=\emptyset$ (3) $V^{def}_{1}\cap V^{use}_{2}=\emptyset$, then the two statements are considered independent from each other and can be swapped. For instance, in the original program shown in Fig. 3, statements on line 4 and line 5 constitute a pair of independent statements. Swapping such pairs of statements will not disrupt the original data flow and is semantic-preserving.
  • Name Random. This perturbation assigns randomly generated names to all program variables and methods. Consistent with Def-use Break, the generated names consist of five random letters or digits. Altering the variable name of a program does not affect program semantics, given that the altered variable names do not conflict with others. The corresponding variable names in the specifications also need to be changed accordingly.
  • Name Shuffle. The idea of this perturbation is the same with Name random but is done by reshuffling all the variable names within the original program rather than generating new names. The corresponding specifications are changed accordingly. Similarly, the perturbation does not affect the original program semantics.
It is worth noting that for some of the perturbations, the specifications involved in the original program should be mutated simultaneously. For instance, Name Random and Name Shuffle alter all the variable names in the original program, where the specifications should be mutated following the same pattern since the variable names are involved in the specifications as well. If-else Flip perturbation may change the positions of some specifications within the target if-else branch, so the specifications involved need to be migrated. Such modifications to specifications pose no impact on their semantics. Additionally, task prompts involving specifications have to be mutated as well, keeping them equivalent to the original tasks, such as the candidate specifications used in the prompt in Specification Correctness Judgement and Specification Candidate Selection tasks.
3.3.2. Evaluation Metrics
We adopt different metrics for different tasks for counterfactual analysis. For the task of Judgement, Selection, and Infilling, we adopt Jaccard Distance to measure the impact of perturbations on the task as $J=1-|\mathbb{P}\cap\mathbb{P}^{\prime}|/|\mathbb{P}\cup\mathbb{P}^{\prime}|$ , where $\mathbb{P}$ and $\mathbb{P^{\prime}}$ refer respectively to the set of original and perturbed programs that an LLM can successfully handle for one kind of perturbation in Section 3.3.1. The programs in $\mathbb{P}\cap\mathbb{P}^{\prime}$ refer to the LLM successfully handles both the original and the perturbed versions, indicating that its performance is consistent on these programs, without being interfered by the perturbation, whereas other programs in $\mathbb{P}\cup\mathbb{P}^{\prime}$ witness a varied performance on different versions. The Jaccard Distance measures the impact of perturbations by calculating the percentage of inconsistently-performing programs. A high distance indicates a strong impact that the corresponding perturbation poses.
For the task of Generation, since it is quantitatively measured by precision and recall, we define Average Variance to measure the impact of perturbations for this task. For a specific type of perturbation in Section 3.3.1, the Average Variance of metric $m$ where $m\in\{Precision,Recall\}$ can be calculated as follows:
$v_{m}=\frac{1}{|\mathbb{P}|}\sum\limits_{(P,P^{\prime})\in\mathbb{P}}|m(P^{\prime})-m(P)|\ \quad where\quad\ \mathbb{P}=\{(P,P^{\prime})\ |\ m(P)>0\vee m(P^{\prime})>0\}$
$(P,P^{\prime})$ denotes pairs of the original program and the perturbed program. Note that we only consider $(P,P^{\prime})$ where either $P$ or $P^{\prime}$ achieve a non-zero value of $m$. The average variance of precision and recall is denoted as $v_{prec}$ and $v_{rec}$, respectively.
3.4. Progressive Consistency Analysis
Progressive Consistency describes the degree to which a model can sustain its performance across a sequence of interrelated tasks. (Chen et al., 2024) Intuitively, for a series of sequentially related tasks, models that fail with the preliminary tasks should tend to fail the advanced follow-up tasks as well. The essence of progressive consistency lies in the sequential dependencies within the series of tasks. The completion of follow-up tasks should partially rely on the knowledge obtained from previous tasks or the ability to complete previous tasks. Given a series of tasks designed with the criterion above, we can evaluate the performance consistency between these tasks to assess whether the models comprehend the code progressively.
3.4.1. Sequential Dependencies Between Evaluation Tasks
Concerning the four evaluation tasks described in Section 3.2, the sequential dependency in between mainly lies in the ability to complete them. Each follow-up task relies partially on the ability to complete its previous task.
  • From Judgement to Selection. The selection task essentially requires models to judge the correctness and strength of all given candidates respectively and choose the correct and strongest one based on the judgments.
  • From Selection to Infilling. To complete an infilling task, the models have to comprehend the specification to fill in, analyze the semantics of the missing parts, select the most appropriate code patterns in the program, and convert them to formalized specifications.
  • From Infilling to Generation. Essentially, the specification generation tasks require multiple rounds of infilling without additional contexts for completion. Generating strong specifications requires a thorough understanding of LLMs concerning the semantics of the target program.
With the intuitions above, we can confirm the sequential dependency existing between the evaluation tasks in SpecEval.
3.4.2. Evaluation Metrics
To quantitatively evaluate the consistency described above, we design a specialized metric, Progressive Consistency Score, or PCS for short, for evaluation. The core idea of PCS is to measure how many tasks in a row the model can handle. Due to the sequential dependency described in Section 3.4.1, we expect the models to handle tasks consecutively given that their semantics learning abilities towards sequential tasks with progressive difficulties are leveraged in a progressive manner as well. We denote the results of the four tasks concerning the same program $P$ as a sequence $R(P)=[r_{1},r_{2},r_{3},r_{4}]$, where $r_{i}\in\{\mathbf{true},\mathbf{false}\}$ denotes whether the model succeeds in the $i$-th task. For Judgement, Selection, and Infilling, the tasks are considered successful if the answer equals the ground truth or passes verification. For Generation, we adopt softened criteria and view those trials with $Precision(P)>0$ as successful generations since positive precision indicates some correct specifications (that can pass the verifier) generated and partial semantics comprehended by LLM. Specifically, the PCS for program $P$ (denoted as $\textbf{PCS}(P)$) can be calculated as follows:
  • If the model successfully handles all four tasks, i.e., $R(P)=[\mathbf{true},\mathbf{true},\mathbf{true},\mathbf{true}]$, it receives a full score of $\textbf{PCS}(P)=1$.
  • If the model handles three tasks in a row, i.e., $R(P)=[\mathbf{true},\mathbf{true},\mathbf{true},\mathbf{false}]$ or
    $R(P)=[\mathbf{false},\mathbf{true},\mathbf{true},\mathbf{true}]$, it receives a half score of $\textbf{PCS}(P)=0.5$.
  • If the model handles two tasks in a row, for instance, $R(P)=[\mathbf{true},\mathbf{true},\mathbf{false},\mathbf{false}]$ or
    $R(P)=[\mathbf{true},\mathbf{false},\mathbf{true},\mathbf{true}]$, it receives a quarter score of $\textbf{PCS}(P)=0.25$.
  • If the model can only handle no consecutive tasks such as $R(P)=[\mathbf{true},\mathbf{false},\mathbf{false},\mathbf{true}]$ or fails to handle any task, we recognize that no consistency is shown between the tasks, and it receives a score of $\textbf{PCS}(P)=0$.
Correspondingly, the average Progressive Consistency Score for the set $\mathbb{P}$ of all programs in the benchmark can be calculated as $\mathbf{PCS}=\frac{1}{|\mathbb{P}|}\sum_{P\in\mathbb{P}}\mathbf{PCS}(P)$. It can be observed that PCS is defined based on the number of consecutive tasks handled by the model. Higher scores indicate more tasks consecutively handled and intuitively higher progressive consistency in the model’s performance. Additionally, considering the difficulty of completing all four tasks, the score obtained is designed to grow exponentially from 0.25 to 1 as the number of consecutively handled tasks increases. Compared to the Incremental Consistency Score (ICS) designed by Chen et al. (Chen et al., 2024) to evaluate the reasoning ability concerning program runtime behaviors, PCS considers more potential consistency between all the tasks, especially the follow-up ones. For instance, $R(P)=[\mathbf{false},\mathbf{true},\mathbf{true},\mathbf{true}]$ receives an ICS of 0, whereas the score is 0.5 for PCS, since such performance still demonstrates consistency between the follow-up tasks. We assume that the consistency between the latter tasks is meaningful despite the failure in preliminary tasks and relevant inconsistency.
4. experimental setup
4.1. Benchmark Construction
We construct a benchmark that includes programs and corresponding ground-truth specifications for evaluation. Concerning the specification language, we adopt the Java Modeling Language (JML) (Leavens et al., 1998) for Java programs for evaluation since they can be conveniently validated by OpenJML (Cok, 2011). We use the released datasets by the latest two research of LLM-based specification generation (Wen et al., 2024; Ma et al., 2024) for evaluation. One work by Ma et al. (Ma et al., 2024) utilizes two datasets, SpecGenBench and the Java Category of SV-COMP Benchmark (sosy lab, 2024). The former contains 120 Java programs with manually-written strong specifications and the latter is widely adopted in the evaluation of software verification tools. For SpecGenBench, we extract 117 programs from the dataset along with corresponding specifications, whereas 3 programs are discarded due to incompatibility with OpenJML. For the Java Category of SV-COMP Benchmark, as these programs are without corresponding specifications, we manually craft them following the similar procedure adopted in the construction of SpecGenBench (Ma et al., 2024). Extensive efforts have been conducted to manually write these specifications, and we have succeeded in 47 programs from SV-COMP that are compatible with OpenJML, and the written specifications can fully specify the program semantics. Another work (Wen et al., 2024) leverages another dataset, Frama-C-Problems (github, 2024), which contains 47 C programs and corresponding ACSL (Baudin et al., 2008) specifications. To cater to Java style, we manually rewrote the programs in Java and altered the specifications to JML-style. 40 programs that can be equivalently translated into Java programs. Others are discarded due to specific features (i.e., pointers) of the C programming language that cannot be trivially translated.
All specifications have been cross-validated by three experts in formal methods and programming, where the semantics of the specifications and corresponding programs are compared. The specifications will be re-formulated until their strength is confirmed by all the experts. This is to ensure high quality within the specifications and ability to describe the behaviors and functionalities of the target program thoroughly. Additionally, all specifications are also ensured to be verifiable, i.e., able to pass the verification of the verifier (namely OpenJML (Cok, 2011) in our work), so the correctness of the specifications can be guaranteed. Eventually, 204 programs are collected for evaluation.
During the experiment, we adopt dynamic validation supported by specification verifiers to check the correctness of the specifications generated by subject LLMs (i.e., to calculate $\mathbf{Cr}(s)$ mentioned in Section 3.2 for a given specification $s$). A specification is considered correct (i.e., $\mathbf{Cr}(s)=\mathbf{true}$) if it maintains its consistency with the target program during a runtime execution process, where test cases are utilized to invoke the target programs. To this end, test cases are also prepared for each program. Similar to the preparation of ground-truth specifications, we inherit existing test cases in the original programs if there are or employ experts to manually craft them otherwise. The experts are required to cover as many branches as possible with the written test cases. False specifications (those in Section 3.2.1 and Section 3.2.2) are generated to test whether the test cases can disprove them, and more test cases should be crafted if they failed. Each program receives, on average, 15.10 test cases, with an average branch coverage of 95.54% and line coverage of 93.34%. The average LoC (Lines of Code) and CC (Cyclomatic Complexity) of programs in the benchmark are 20.06 and 6.34, respectively. Detailed statistics of the benchmark are provided on our website (Author, s).
4.2. Studied LLMs
In this work, we selected six state-of-the-art and popular LLMs for evaluation.
  • GPT-3.5-Turbo (OpenAI, 2023), a general LLM developed by OpenAI (OpenAI, 2024a), able to understand and generate natural language or code. We adopt gpt-3.5-turbo-0125, the latest model in the GPT-3.5 family during the experiment, for evaluation.
  • GPT-4-Turbo (Achiam et al., 2023), the latest general LLM developed by OpenAI, with the most state-of-the-art performance among all general LLMs currently.
  • Llama 2 (Touvron et al., 2023), an open-source general LLM developed by Meta AI (AI, 2024). We adopt Llama-2-7b-chat for evaluation.
  • CodeLlama (Roziere et al., 2023), an open-source code LLM based on the architecture of Llama 2. We adopt CodeLlama-7b-Instruct for evaluation.
  • Deepseek-Coder (Guo et al., 2024), an open-source code LLM employing the same architecture as DeepSeek LLM (Bi et al., 2024). We adopt the version deepseek-coder-6.7b-instruct for evaluation.
  • Magicoder (Wei et al., 2023), an open-source code LLM fine-tuned with OSS-Instruct, a novel high-quality code instruction data. The version employed in the evaluation is Magicoder-S-DS-6.7B.
The selected models have been applied to a wide variety of code-related tasks in existing works (Chen et al., 2021; Liu et al., 2023; Deng et al., 2023). They also exhibit diversities over multiple dimensions, ranging from general LLMs to code LLMs, open-source models to closed-source models, and foundation models to fine-tuned models. Due to the computing resource limit, only 7B versions of open-source LLMs are selected.
4.3. Prompt Format and Model Configurations
In terms of the prompts provided to LLMs, we generally adopt a unified prompt structure for each task and all the subject LLMs in the experiments. The structure mainly consists of four components. Initially, a system message is appended to the context to inform the model of its basic role. The few-shot prompting technique is then adopted, where several pairs of example requests and replies are provided to the LLM, assisting the LLM in learning the task requirements and desired output format. Afterward, a description of the target task (i.e., the current task that the LLM is expected to answer) is appended to the context. Up to this point, the context constitutes a complete query and is sent to the LLM, which learns necessary information from the context and completes the task within its reply. The template prompts for different tasks are provided on our website (Author, s).
For the open-source models mentioned in Section 4.2, we deploy a local server to execute the models and process relevant queries, running on a Linux Server equipped with one H800 GPU with 80GB memory. The source code and model weights of all open-source models can be accessed at corresponding HuggingFace (HuggingFace, 2024) repositories. For the closed-source models (i.e., GPT-3.5 and GPT-4), we access them through the API provided by OpenAI (OpenAI, 2024b). Following the settings of previous works (Du et al., 2024; Liu et al., 2024a), we adopt greedy decoding (Chen et al., 2015) for all subject LLMs to make the results reproducible. Generally, the models are configured as top_k=1 and temperature=0 under greedy decoding. The max context length is set to 2048 tokens, leaving abundant space for the LLMs to generate detailed results. Other parameters are kept consistent with default model settings.
5. Experimental Results
We aim to answer the following research questions through evaluation:
  • RQ1: How is the performance of LLMs on specification-related tasks generally? This gives us an overall scenario for the performance of LLMs on each task.
  • RQ2: How is the performance of LLMs against semantic-preserving perturbations? This evaluates the impact of the perturbations through Counterfactual Analysis.
  • RQ3: Can LLMs maintain performance consistency over progressive tasks? It assesses the consistency of LLMs over different tasks through Progressive Consistency Analysis.
5.1. RQ1: Overall Performance
Table 1 shows the overall performance of each model on four tasks for all versions of programs (including original and five perturbed variants).
5.1.1. Comparison Between Tasks
For different tasks involved in the experiments, the models showcase different levels of fulfillment. We take the original category as the example for illustration. For the Candidate Judgement task, the majority of models can achieve an accuracy above 50%, with some excelling models achieving relatively high accuracy, such as the 81.37% of GPT-4. For the Candidate Selection task, the accuracy of most models shows some decline compared to that of Judgement (from 57.84% to 29.41% of CodeLlama) or has no significant variance (from 44.12% to 39.71% of Magicoder). When it comes to the Specification Infilling task, we can witness a significant drop in the performance of all models among all categories, ranging from 7.84% of Llama to 31.86% of Magicoder. As for the Specification Generation task, the average precision of most models ranges from 24.01% of Llama to 54.05% of GPT-3.5, whereas the average recall is relatively lower, varying between 15.88% of Llama and 37.80% of GPT-3.5. The low recall of all LLMs suggests that they can only generate limited specifications with strength comparable to the ground truth, indicating a relatively weak ability to summarize full program semantics into the form of formal specifications. The tasks for the other five categories showcase similar performance variance patterns as well.
According to their performance, the most straightforward task involved in the experiment is Judgement, followed by Selection, then Generation, whereas the most challenging task is Infilling. It is an interesting finding, and to our intuition, the root cause of this phenomenon lies in the different aspects of ability that are examined by different tasks. Judgement and Selection are mainly about reading and understanding (i.e., ”input”) specifications, where LLMs try to comprehend given specifications and give relevant conclusions. On the contrary, Generation is about writing and expressing (i.e., ”output”) specifications, where LLMs can freely articulate code semantics they understand in the form of specifications. The criteria for successful generation is to pass the verifier. Thus, the written specifications may not be strong. The tasks mentioned above examine two opposite aspects of ability concerning specifications. Interestingly, Infilling lies somewhere in between, where LLMs write partial specifications but cannot do it at any will. They have to generate results within the structure defined by the target specification to fill in, which are usually strong with sufficient semantic expression. To this end, they have to read and understand the given specification first; some may be difficult for them, and then, based on their understanding, write the standard answers. Consequently, the Infilling task examines both of the two aspects mentioned above (i.e., ”input” and ”output”) simultaneously, making it the most difficult task of all.
Table 1. Performance of all models on each task and perturbation category. Acc.: Average Accuracy, AvgPrec.: Average Precision, AvgRec: Average Recall in percentages. #Pass: the number of programs handled with all test cases passed.
Category LLM Judgement Selection Infilling Generation Category LLM Judgement Selection Infilling Generation
Acc. Acc. Acc. #Pass AvgPrec. AvgRec. #Pass Acc. Acc. Acc. #Pass AvgPrec. AvgRec. #Pass
Original GPT-3.5 64.71 67.65 23.53 48 54.05 37.80 79 Def-use Break GPT-3.5 67.65 61.27 27.45 56 53.37 38.47 77
GPT-4 81.37 89.71 11.27 23 32.90 22.75 54 GPT-4 78.92 86.27 14.22 29 24.28 16.91 34
Llama 50.00 30.88 7.84 16 24.01 15.88 21 Llama 54.41 24.02 5.39 11 20.98 16.02 16
CodeLlama 57.84 29.41 17.65 36 43.70 26.05 61 CodeLlama 62.25 28.92 15.69 32 35.32 23.04 43
Deepseek-coder 57.35 61.27 17.65 36 43.38 31.37 55 Deepseek-coder 60.29 59.80 10.29 21 40.79 29.99 46
Magicoder 44.12 39.71 31.86 65 53.22 34.95 67 Magicoder 53.43 45.59 31.86 65 48.44 35.44 54
If-else Flip GPT-3.5 68.14 65.20 22.55 46 42.61 31.26 62 Independent Swap GPT-3.5 66.18 66.18 25.00 51 51.53 39.02 75
GPT-4 77.94 83.82 11.76 24 20.20 14.20 28 GPT-4 79.90 86.27 12.75 26 23.78 18.23 32
Llama 49.51 24.51 8.82 18 28.59 21.71 23 Llama 48.53 22.55 7.84 16 24.53 17.82 20
CodeLlama 63.24 30.88 13.24 27 38.28 27.79 43 CodeLlama 65.69 35.29 14.71 30 36.26 25.38 42
Deepseek-coder 61.27 57.35 17.16 35 41.79 30.93 55 Deepseek-coder 55.88 59.80 18.14 37 42.07 31.03 55
Magicoder 55.88 50.00 32.35 66 47.33 35.95 53 Magicoder 58.82 43.14 35.29 72 44.35 32.85 52
Name Random GPT-3.5 63.24 54.90 24.02 49 48.49 37.16 74 Name Shuffle GPT-3.5 68.14 66.67 20.59 42 41.43 30.99 57
GPT-4 76.96 81.37 10.78 22 25.55 18.57 40 GPT-4 76.96 84.31 12.75 26 29.52 21.26 46
Llama 56.37 25.00 10.78 22 22.79 16.45 20 Llama 49.02 26.96 7.35 15 28.88 19.77 31
CodeLlama 59.31 25.00 15.20 31 38.35 25.00 53 CodeLlama 59.80 30.39 15.69 32 33.47 24.48 37
Deepseek-coder 55.88 61.27 18.63 38 37.54 28.32 45 Deepseek-coder 59.80 59.31 13.73 28 36.75 28.63 41
Magicoder 60.29 48.04 36.76 75 43.79 34.73 54 Magicoder 63.73 46.08 28.43 58 45.57 35.03 51
5.1.2. Comparison Between Models
It can be observed that different models are specialized in different tasks. Regarding Judgement, GPT-4 outperforms all other models significantly with leading accuracy across all categories, followed by GPT-3.5, of which the accuracy is ranked second in all categories. The three code LLMs involved in the experiments underperform the GPT family on this task. Regarding Selection, a similar pattern can be observed across all categories, where GPT-4 is way ahead, followed by GPT-3.5, then the three code LLMs. It is also worth noting that Llama and CodeLlama, two models with the same architecture, suffer from significant performance declines in Selection tasks compared to Judgement among all categories. As described in Section 5.1.1, Judgement and Selection are mainly about reading and understanding specifications. This indicates that GPT-4 and GPT-3.5 showcase impressive ability regarding the comprehension of specifications.
Nevertheless, the ranking varies significantly when it comes to the latter two tasks. In terms of Infilling, the best-performing model is Magicoder, the only model achieving an accuracy above 30% across 5 out of 6 categories. The following model is GPT-3.5, another model that can keep an accuracy above 20% among all categories. Other models demonstrate a relatively lower performance, with accuracy below 20% in all categories. In terms of Generation, the leading model is GPT-3.5, with the highest average precision and recall over 4 out of 6 categories, along with the highest #Pass over all 6 categories. The second-best model is Magicoder, able to surpass GPT-3.5 in 2 out of 6 categories concerning precision and recall.
Thus, we can find that, despite the excelling performance of the closed-source GPT models by OpenAI, it is still possible for open-source models like Magicoder to achieve comparable or even better performance with a much smaller model scale under the assistance of fine-tuning with high-quality code data corpora.
Answer to RQ1: LLMs showcase a limited ability to articulate program semantics with formal specifications. The best-performing task is Judgement, followed by Selection, then Generation, and the most challenging task is Infilling. Notably, Magicoder exhibit leading performance in Infilling and Generation, indicating the possibility for small-scale LLMs to surpass large-scale ones.
Table 2. Impact of all types of perturbations concerning different models and tasks. DUB: Def-use Break. IEF: If-else Flip. IDS: Independent Swap. NMR: Name Random. NMS: Name Shuffle. $J_{jud}$: Jaccard Distance for Judgement. $J_{sel}$: Jaccard Distance for Selection. $J_{inf}$: Jaccard Distance for Infilling. $v_{prec}$: Average Variance of Precision. $v_{rec}$: Average Variance of Recall.
Model Metric DUB IEF IDS NMR NMS Overall Model Metric DUB IEF IDS NMR NMS Overall
GPT-4 $J_{jud}$ 0.153 0.110 0.065 0.144 0.165 0.455 GPT-3.5 $J_{jud}$ 0.323 0.337 0.362 0.358 0.348 0.419
$J_{sel}$ 0.120 0.107 0.101 0.163 0.141 $J_{sel}$ 0.303 0.263 0.239 0.366 0.277
$J_{inf}$ 0.818 0.825 0.860 0.714 0.805 $J_{inf}$ 0.535 0.484 0.586 0.594 0.657
$v_{prec}$ 0.630 0.699 0.775 0.623 0.719 $v_{prec}$ 0.490 0.499 0.445 0.495 0.527
$v_{rec}$ 0.518 0.523 0.586 0.460 0.539 $v_{rec}$ 0.399 0.399 0.345 0.422 0.422
Avg. 0.448 0.453 0.477 0.421 0.474 Avg. 0.410 0.396 0.395 0.447 0.446
Llama $J_{jud}$ 0.457 0.496 0.511 0.534 0.493 0.609 CodeLlama $J_{jud}$ 0.275 0.248 0.250 0.340 0.235 0.409
$J_{sel}$ 0.845 0.798 0.840 0.837 0.854 $J_{sel}$ 0.413 0.292 0.308 0.458 0.373
$J_{inf}$ 0.962 0.828 0.667 0.848 0.852 $J_{inf}$ 0.612 0.535 0.500 0.759 0.583
$v_{prec}$ 0.536 0.527 0.533 0.538 0.542 $v_{prec}$ 0.500 0.447 0.488 0.455 0.502
$v_{rec}$ 0.346 0.364 0.349 0.338 0.337 $v_{rec}$ 0.326 0.339 0.325 0.314 0.345
Avg. 0.629 0.603 0.580 0.619 0.616 Avg. 0.425 0.372 0.374 0.465 0.408
Deepseek -Coder $J_{jud}$ 0.389 0.376 0.362 0.396 0.374 0.441 Magicoder $J_{jud}$ 0.526 0.563 0.521 0.521 0.523 0.485
$J_{sel}$ 0.236 0.284 0.210 0.252 0.230 $J_{sel}$ 0.597 0.524 0.580 0.579 0.529
$J_{inf}$ 0.643 0.709 0.596 0.679 0.720 $J_{inf}$ 0.632 0.511 0.425 0.571 0.570
$v_{prec}$ 0.517 0.506 0.488 0.492 0.501 $v_{prec}$ 0.434 0.418 0.452 0.463 0.450
$v_{rec}$ 0.399 0.426 0.413 0.402 0.422 $v_{rec}$ 0.319 0.348 0.345 0.355 0.371
Avg. 0.437 0.460 0.414 0.444 0.449 Avg. 0.502 0.473 0.465 0.498 0.489
5.2. RQ2: Counterfactual Analysis
The Jaccard Distances and Average Variances of different models on different tasks are listed in Table 2. As described in Section 3.3, both denote the perturbations’ impact. A Higher Jaccard Distance or Average Variance indicates a higher impact imposed by the perturbation. Generally, all models exhibit performance variance to different degrees, indicating that the semantics LLMs learn are interfered with by the perturbations. It can be observed that the impact of perturbations varies differently among different tasks, and the impact tends to be stronger in the tasks where the models underperform. For instance, as described in Section 5.1, GPT-4 significantly outperforms all the other models in terms of Judgement and Selection task, where GPT-4 also possesses the lowest $J_{jud}$ and $J_{sel}$, showing performance consistency over all types of perturbations on these two tasks. Also, the performance of all models showcases a significant decline in the Infilling tasks, and correspondingly, the $J_{inf}$ metrics of all models are also significantly higher than that of other tasks. The models that perform relatively better in Infilling, such as Magicoder and GPT-3.5, also showcase stronger resistance against the perturbations with relatively lower $J_{inf}$.
We further measure the average impact of different perturbations on each model across 5 tasks. It is calculated by the average of all impact metrics of each perturbation across 5 tasks. The Avg. row presents the average impact of different perturbations in Table 2. We can observe that different models are sensitive to different types of perturbations. For instance, GPT-4 is sensitive to Independent Swap (i.e., IDS in Table 2) and Name shuffle (i.e., NMS), with an exceptionally high variance of 0.477 and 0.474 in these categories. GPT-3.5 is relatively sensitive to Name Random (i.e., NMR) and Name shuffle, Deepseek-Coder is relatively weak at If-else Flip (i.e., IEF), and Magicoder is relatively weak at Def-use Break (i.e., DUB). Concerning the average impact on all models, Name Random achieves the highest average impact of 0.482, followed by Name Shuffle, with an average impact of 0.480. These two types of perturbations showcase the most decisive influence on model outputs. The following perturbation is Def-use Break, with a slightly lower impact of 0.475. The other two types of perturbation, If-else Flip and Independent Swap, exhibit a relatively lower impact of 0.460 and 0.451, respectively.
We also calculated the overall impact of all perturbations on each model (presented in the columns starting with Overall in Table 2). Across all subject LLMs, CodeLlama showcases minimal performance variance against all types of perturbations, with the average impact on it being 0.409. The following model is GPT-3.5, with a slightly higher performance variance of 0.419. In general, the two leading models exhibit better resistance against perturbations. Other models demonstrate higher performance variance and more sensitivity to the perturbations.
Answer to RQ2: All LLMs exhibit sensitivities of different levels towards the perturbations, indicating that perturbations can interfere with the semantics learned. The impact of perturbations varies differently among different tasks, with a tendency to be stronger in the tasks where the models underperform.
Table 3. Progressive Consistency Score of each model on all perturbation categories.
Model Original Def-use Break If-else Flip Independent Swap Name Random Name Shuffle
GPT-3.5 0.276 0.272 0.263 0.266 0.240 0.237
GPT-4 0.262 0.260 0.229 0.248 0.238 0.243
Llama 0.124 0.121 0.127 0.120 0.135 0.124
CodeLlama 0.169 0.174 0.159 0.182 0.159 0.169
Deepseek-Coder 0.218 0.193 0.214 0.223 0.208 0.206
Magicoder 0.214 0.237 0.249 0.246 0.257 0.238
5.3. RQ3: Progressive Consistency Analysis
The Progressive Consistency Scores (PCS) of each model on all perturbation categories are presented in Table 3. Generally, all LLMs achieve a relatively low PCS under 0.3, indicating that all LLMs exhibit limited consistency between the tasks involved in the experiments. Considering the relatively low performance of all subject LLMs on the third task Infilling mentioned in Section 5.1, this phenomenon is reasonable. Underperformance on the third task indicates that LLMs can only solve at most two tasks in a row for the majority of programs. The model exhibiting the highest consistency is GPT-3.5, with an average PCS above 0.25 among 4 out of 6 categories. The following model is GPT-4, with a small step behind GPT-3.5 in terms of PCS. The PCS of the above two models is around 0.25, indicating that most of the programs succeed in handling two tasks consecutively. Deepseek-Coder and Magicoder form the second tier, both achieving an average PCS above 0.2 in most categories. Llama and CodeLlama achieve a relatively lower PCS below 0.2, exhibiting the least consistency among all models.
This observation shows that code LLMs do not necessarily exhibit better progressive consistency than general LLMs, despite their performance on each task mentioned in Section 5.1. Since GPT-4 and GPT-3.5 are closed-source commercial LLMs trained with larger model scale and possibly more immense data corpora, they exhibit better text comprehension and logical reasoning abilities, which can be leveraged in our tasks, especially the first task Judgement and the second task Selection. This leads to the solid and excelling performance of these two models on the former two tasks, guaranteeing a stable PCS above 0.25. Despite the impressive abilities showcased by open-source models such as Magicoder in the latter two tasks, Infilling and Generation, the tasks are relatively more challenging and cannot be easily handled in a row. Limited Progressive Consistency Scores can be earned through them, eventually leading to the gap behind GPT-4 and GPT-3.5.
In terms of consistency across different categories, we can witness declines within 3~4 out of 6 models for each perturbed category compared to Original category, but usually not significant. Nevertheless, in the perturbations with the strongest impact, i.e., Name Random and Name Shuffle, we can witness some decline in the consistency of GPT-3.5 and GPT-4. However, Magicoder achieves its highest PCS in Name Random across all categories, indicating that the impact of perturbations on the performance consistency varies across different models.
Answer to RQ3: The LLMs exhibit limited progressive consistency, indicating that they are not reasoning in a progressive manner (top-down or bottom-up) as humans do. GPT-4 and GPT-3.5 showcase a relatively better consistency due to their excelling performance in Judgement and Selection tasks. Future research remains focused on enhancing the performance of LLMs in tasks involving continuous and contextually linked reasoning.
6. discussion
6.1. Implication
The potential implications of this work cover a list of areas. Concerning the Evaluations on LLMs, semantic-preserving perturbations are proved to have a non-negligible impact on the performance of LLMs, which is worth attention when designing further benchmarks or evaluation frameworks. The resistance against perturbations should also be emphasized when developing new language models. Regarding Code Comprehension with LLMs, the assistance of program specifications in natural language or formal languages deserves to be noted. In terms of Code Completion and Generation with LLMs, the possibility can be explored concerning the utilization of specifications to assist the quality assurance of LLM-generated code. Concerning the Software Development process, the importance of all types of specifications (including documents and comments) should not be overlooked since they convey rich information about code and improve its readability and maintainability. As for Education, it is possible that program specifications in natural or formal languages can assist students in understanding example code during the teaching process.
6.2. Greedy Decoding for Model Generation
Following previous works (Liu et al., 2023; Du et al., 2024; Liu et al., 2024a), we adopted greedy decoding (top_k=1 and temperature=0) for all subject LLMs in the experiments. This is to guarantee the determinism in the sampling process of LLMs and the reproducibility of the final results. Nevertheless, some research within the same domain (Gu et al., 2024; Chen et al., 2024) adopted random sampling for evaluation, which is also a choice. Currently, research has emerged to compare the performance of LLMs under greedy decoding and other decoding strategies. Cobbe et al. (Cobbe et al., 2021) and Hendrycks et al. (Hendrycks et al., 2021) show that greedy decoding is preferred when math reasoning problems are to solve. Song et al. (Song et al., 2024) claim that greedy decoding is generally more effective for most tasks, especially reasoning tasks and coding problems. Since the specification-related tasks involved in this work are generally about reasoning and formulating formal statements, greedy decoding tends to be the more suitable choice.
6.3. Threats to Validity
Internal Threats. First, experimental results may be affected by the prompts utilized to communicate with LLMs. We follow the prompt structure of Chen et al. (Chen et al., 2024) when designing the prompts. The effect of prompt strategies on the model performance will be studied in future works. Second, the base datasets, based on which we formulate the adapted benchmark, face potential data leakage when using the LLMs for evaluation. For programs in Frama-C-Problems (github, 2024) and SV-COMP (sosy lab, 2024), the ground-truth JML specifications involved are manually crafted by experts without risks of data leakage. As for SpecGenBench (Ma et al., 2024), the dataset was publicly released in March 2024, whereas the versions of models we adopted were all released prior to this point in time. Thus, using our selected LLMs for evaluation is without data leakage issues. However, if the selected models are after March 2024, there may be potential data leakage risks if the data is used for training.
External Threats. The potential threat lies in the specification verifier. OpenJML (Cok, 2011) provides static verification and runtime checking for validation. Static verification requires sufficient supporting information (i.e., other strong specifications) and fails easily if not given enough, introducing serious bias. Therefore, runtime checking, with more stability and reliability, is adopted instead.
7. conclusion
In this paper, we present SpecEval, a novel black-box framework evaluating the code comprehension ability in Large Language Models via program specifications. Leveraging the feature that specifications cover abundant execution traces and rich semantic information, we adopt them as an effective representation of program semantics. Four specification-related tasks with sequential dependencies are designed to assess the performance of LLMs. Counterfactual Analysis and Progressive Consistency Analysis are adopted to analyze the performance consistency of LLMs over different dimensions. Experimental results show that LLMs showcase a limited ability to fully articulate program semantics with formal specifications. Counterfactual Analysis and Progressive Consistency Analysis also reveal the inconsistency in their performance over different dimensions.